OpenStudy (anonymous):

cos^2 - sin^2X ------over------ 1 - tan^2 X

Nnesha (nnesha):

\[\huge\rm \frac{ \cos^2 x- \sin^2x }{ 1-\tan^2x}\]

OpenStudy (anonymous):

is cos^2X the answer

Nnesha (nnesha):

can you please tell me how u got cos^2x :=) thanks!

OpenStudy (anonymous):

do i have to

Nnesha (nnesha):

hmm just wanted some explanation you can explain in words like what ws ur first step

OpenStudy (anonymous):

cos^2 x - sin^2 x cos^2 x -sin^2 x cos^2 x - sin^2 x -------------------- = --------------------- = ---------------------- 1 - tan^2 cos^2 x sin^2 x cos^2 x - sin^2 x -------- - --------- ----------------------- cos^2 cos^2 cos^2 x cos^2 - x - sin^2 cos^2 x = ------------------------- x -------------- = cos^2 x 1 cos^2 -sin^2 x

Nnesha (nnesha):

PERFECT!

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

thanks for showing ur work and you can use equation tool great job!

OpenStudy (anonymous):

yea i'm new so i don't really know how to use this yet

Nnesha (nnesha):

yeh i understand let me know if you have any question abt equation tool i can help :=)

Nnesha (nnesha):

or about latex

OpenStudy (anonymous):

ok thanks

Nnesha (nnesha):

np :=)